# JETSCAPE Collaboration

### v1

@@ III Analyses
This section provides a detailed description of the Soft Drop grooming procedure used for calculating observables that characterize the hard component structures of jets, as well as the method for constructing $\gamma$-tagged jets in our Monte Carlo simulations of high-energy heavy-ion collisions.
III.1 Soft Drop grooming procedure
In this study, we focus on observables characterizing hard splittings of jets obtained through the Soft Drop grooming algorithm [62]. The Soft Drop grooming identifies the hard splitting, while removing the soft branchings, at as large an angle as possible.
In the Soft Drop grooming procedure, an angular-ordered clustering tree is rebuilt using the Cambridge-Aachen (C/A) algorithm [133, 134] for constituents of a triggered jet reconstructed by a standard jet-finding algorithm, such as the anti-$k_{t}$ algorithm [135], with a jet cone size $R$. Next, we traverse back, i.e., from the branch with the largest angle, through the C/A tree. At each branching, we check whether the two prongs of the branching satisfy the Soft Drop condition given by
$\frac{p_{T,1},p_{T,2}}{p_{T,1}+p_{T,2}}>z_{\mathrm{cut}}\left(\frac{\Delta R_{12}}{R}\right)^{\beta},$ (1)
where $p_{T,1}$ and $p_{T,2}$ ($<p_{T,1}$) are the transverse momenta of the prongs, and $\Delta R_{12}=\sqrt{(\eta_{1}-\eta_{2})^{2}+(\phi_{1}-\phi_{2})^{2}}$ is the radial distance between the prongs in the rapidity-azimuthal angle plane. The Soft Drop parameters $z_{\mathrm{cut}}$ and $\beta$ control the grooming procedure.
If the Soft Drop condition is satisfied, we stop the procedure, and the two prongs of the branching are used to calculate the groomed jet observables. If the condition is not met, we continue traversing the tree by following the prong with the larger $p_{T}$ and repeat the same procedure. It should be noted that there may be cases where a prong pair satisfying the Soft Drop condition cannot be found eventually. Throughout this work, the jet reconstruction and Soft Drop grooming are performed using the fastjet package [136, 137] with fastjet-contrib-1.045 [138].
@@ III.2 $\boldsymbol{\gamma}$-tagged jet
However, experimentally, $\gamma$-jet pairs can also arise from processes other than the initial hard $\gamma$-jet pair creation, such as photon radiation in final state parton showering. Therefore, to suppress contamination from such non-targeted processes, specific cuts are imposed in the experimental analyses: the isolation requirement for photons and the relative azimuth angle cut. To enhance quantitative accuracy in our comparison to experimental data, we impose these cuts on our simulated events.
@@
The relative azimuth angle cut,
$\lvert\phi_{\mathrm{jet}}-\phi_{\gamma}\rvert>\Delta\phi_{\mathrm{cut}},$ (4)
@@ IV Results
This section systematically investigates the modification of the Soft Drop substructures from simulations with simplified initial events generated by the pgun module. A single quark or gluon is generated at one point in the medium and allowed to propagate outward. Strict control over the flavor and energy of the parton allows for a detailed investigation of the flavor and energy dependence of specific substructure observables.
@@ IV.1 Jet splitting momentum fraction
where $p_{T,1}$ and $p_{T,2}$ ($<p_{T,1}$) are transverse momenta of the pair prongs passing the Soft Drop condition.
Figure 1: Distributions of jet splitting momentum fraction $z_{g}$ normalized by the number of all triggered jets (upper panel) and the number of jets passing the Soft Drop condition (lower panel) for the leading jets in events generated with pgun. The jet shower evolution is performed by vacuum matter for the parent parton having a $E_{\mathrm{init}}=140$ GeV. Jets are reconstructed with $R=0.4$ at midrapidity $|\eta_{\mathrm{jet}}|<2.0$. The results are shown for quark jets (solid) and gluon jets (dashed) with different $p^{\mathrm{jet}}_{T}$ triggers, 112, 84, 56, and 28 GeV. The Soft Drop parameters are $z_{\mathrm{cut}}=0.2$ and $\beta=0$.
In Fig. 1, the results for leading jets from pgun simulations for the vacuum case are shown. The initial energy of the parent parton is fixed at $E_{\mathrm{init}}=140$ GeV. In the upper panel, the $z_{g}$ distribution normalized by the number of all triggered jets, including those that did not pass the Soft Drop condition,
@@
is shown, whereas in the lower panel, the one normalized by the number of jets passing the Soft Drop condition,
@@
is shown. Here, $N_{\mathrm{jet}}$ and $N_{\mathrm{SD,jet}}$ are the number of triggered jets and the number of jets passing the Soft Drop condition, respectively.
Gluon jets pass the Soft Drop condition more often than quark jets, and as a result, for the case of the normalization by the number of all triggered jets, the gluon jet consistently exhibits larger values than the quark jet across the entire $z_{g}$ region. However, when normalized by the number of jets passing the Soft Drop condition, the difference between quark jets and gluon jets is nearly negligible.
Figure 2: Ratios of $z_{g}$ distributions for the leading jets in events with the parent parton having a fixed initial energy $E_{\mathrm{init}}=140$ GeV generated by pgun. The jet shower evolution in the QGP medium produced in 0%–10% Pb-Pb collisions at $\sqrt{s_{NN}}=5.02$ TeV is performed by matter$+$lbt. The results are shown for quark jets (solid) and gluon jets (dashed) with different $p^{\mathrm{jet}}_{T}$ triggers, 112, 84, 56, and 28 GeV. The Soft Drop parameters are $z_{\mathrm{cut}}=0.2$ and $\beta=0$.
@@
Next, we delve into the detailed exploration of how the medium effects at high virtuality and low virtuality, respectively, bring about the hard splitting modification. To achieve this, we conduct simulations incorporating only the medium effects of matter, by turning off the medium effects at low virtuality ($Q<2$ GeV), and present the results of these simulations with the pgun initial hard process in Fig. 3. For both quark jets and gluon jets, the modification in the $z_{g}$ distribution is almost invisible, indicating that the modifications observed in the full matter$+$lbt results are governed by the low-virtuality phase, as modeled by the lbt module.
Figure 4: Same as Fig. 1 for jets from hard scatterings in $p$-$p$ collisions at $\sqrt{s}=5.02$ TeV, generated by pythiagun with ISR and MPI. The results are shown for $\gamma$-tagged jets (solid) from prompt photon-generating hard processes (HardQCD:all=off+PromptPhoton:all=on) and inclusive jets (dashed) from inclusive hard processes (HardQCD:all=on+PromptPhoton:all=on) generated at leading order by pythia 8 and with different $p^{\mathrm{jet}}_{\mathrm{T}}$ triggers. For $\gamma$-tagged jets, isolation requirement, relative azimuth angle cut, and the additional cut of $p^{\mathrm{jet}}_{\mathrm{T}}<p^{\gamma}_{\mathrm{T}}$ are imposed.
@@
For the inclusive jet analysis, events with initial scatterings, including all hard QCD 2-to-2 events (HardQCD:all=on+PromptPhoton:all=on) are used. In the distributions normalized by the number of all triggered jets (upper panel), it can be observed that inclusive jets are slightly more likely to pass the Soft Drop condition compared to $\gamma$-triggered jets, due to the larger fraction of gluon jets. When the normalization is done by the number of jets passing the Soft Drop condition, the difference between $\gamma$-triggered jets and inclusive jets disappear, similar to the erasing of differences between gluon and quark jets as seen in the pgun results.
Figure 5: Same as Fig. 2 for jets from hard scatterings at $\sqrt{s_{NN}}=5.02$ TeV, generated by pythiagun with ISR and MPI. The results are shown for $\gamma$-tagged jets (solid) from prompt photon-generating hard processes (HardQCD:all=off+PromptPhoton:all=on) and inclusive jets (dashed) from inclusive hard processes (HardQCD:all=on+PromptPhoton:all=on) generated at leading order by pythia 8 with different $p^{\mathrm{jet}}_{\mathrm{T}}$ triggers. For $\gamma$-tagged jets, isolation requirement, relative azimuth angle cut, and the additional cut of $p^{\mathrm{jet}}_{\mathrm{T}}<p^{\gamma}_{\mathrm{T}}$ are imposed.
Figure 5 shows the medium modification for the $\gamma$-tagged and inclusive jets in Pb-Pb collisions at $\sqrt{s_{NN}}=5.02$ TeV with the same triggers. For both $\gamma$-tagged jets and inclusive jets, modification becomes less significant as $p^{\mathrm{jet}}_{T}$ increases. Especially for jets with $p^{\mathrm{jet}}_{T}>300$ GeV, almost no changes in the shape of the $z_{g}$ distribution are observed. In all $z_{g}$ regions, jets passing the Soft Drop condition are almost uniformly reduced due to medium effects, with the reduction being slightly more pronounced for inclusive jets.
Figure 6: Same as Fig. 2 for $\gamma$-tagged jets from prompt photon-generating hard processes (HardQCD:all=off+PromptPhoton:all=on) generated at leading order by pythia 8 at $\sqrt{s_{NN}}=5.02$ TeV for different $x_{J\gamma}$ ranges. The photon of $160140<p^{\gamma}_{\mathrm{T}}<$ 160 GeV is triggered with isolation requirement, relative azimuth angle cut.
Figure 6 shows our prediction for the $x_{J\gamma}$ dependence in the modification of $z_{g}$ distribution for $\gamma$-tagged jets in Pb-Pb collisions at $\sqrt{s_{NN}}=5.02$ TeV. As seen in the pgun results, there is no significant $p^{\mathrm{jet}}_{T}$-selection bias effect on the $z_{g}$ distribution, and thus, the modification remains almost consistent across all presented $x_{J\gamma}$ ranges.
@@ IV.2 Jet splitting radius
where $\eta_{1}$, $\eta_{2}$, $\phi_{1}$, and $\phi_{2}$ are the rapidities and azimuthal angles of the pair prongs passing the Soft Drop condition.
Figure 7: Distributions of jet splitting momentum radius $r_{g}$ normalized by the number of all triggered jets for the leading jets in events generated with pgun. The jet shower evolution is performed by vacuum matter for the parent parton having a $E_{\mathrm{init}}=140$ GeV. Jets are reconstructed with $R=0.4$ at midrapidity $|\eta_{\mathrm{jet}}|<2.0$. The results are shown for quark jets (solid) and gluon jets (dashed) with different $p^{\mathrm{jet}}_{T}$ triggers, 112, 84, 56, and 28 GeV. The Soft Drop parameters are $z_{\mathrm{cut}}=0.2$ and $\beta=0$.
@@
for leading gluon jets and quark jets from the vacuum pgun simulations are compared. The gluon jets exhibit much wider distributions with a peak at larger $r_{g}$ values than quark jets. This is due to the fact that gluon jets, having a larger Casimir factor and radiating more, tend to be produced with a larger virtuality compared to quark jets. For the jet cone size $R=0.4$, almost no $p^{\mathrm{jet}}_{T}$ cut dependence can be observed in the vacuum case with fixed $E_{\mathrm{init}}=140$ GeV.
Figure 8: Ratios of $r_{g}$ distributions for the leading jets in events with the parent parton having a fixed initial energy $E_{\mathrm{init}}=140$ GeV generated by pgun. The jet shower evolution in the QGP medium produced in 0%–10% Pb-Pb collisions at $\sqrt{s_{NN}}=5.02$ TeV is performed by matter$+$lbt. All the setups are the same as in Fig. 2.
Figure 8 shows the modification of the $r_{g}$ distribution for the pgun jets by the medium created in central Pb-Pb collisions at $\sqrt{s_{NN}}=5.02$ TeV. For gluon jets, a monotonically decreasing trend with increasing $r_{g}$ is observed in cases with a larger $p^{\mathrm{jet}}_{T}$ cut. Then, as the $p^{\mathrm{jet}}_{T}$ cut value is reduced, the decrease rate becomes more gentle, and almost no modification can be seen for $p^{\mathrm{jet}}_{T}/E_{\mathrm{init}}>0.2$. Thus, for gluon jets, the modification pattern is entirely brought by the effect from jet selection with $p^{\mathrm{jet}}_{T}$ cut than the actual structure modification of the hard splitting: Jets with an originally large splitting radius tend to have their constituents reach the edge of the jet cone more easily, making them lose more energy and thus less likely to be triggered.
The selection bias effect with $p^{\mathrm{jet}}_{T}$ cut can be observed also for quark jets as large-$r_{g}$ suppression, with the same trend disappearing as the $p^{\mathrm{jet}}_{T}$ cut value is reduced. Additionally, in the case of quark jets, a prominent bump structure accompanied by slight suppression at small $r_{g}$ is observed in the mid-$r_{g}$ region. This feature is significantly different from that of gluon jets, providing evidence of direct modifications to the hard-splitting structure within quark jets.
@@
To investigate the hard splitting modification in more detail, we show the matter-only results, in which the medium effect is turned off for $Q<2$ GeV, in Fig. 9. For both quark jets and gluon jets, $r_{g}$ distribution modification for any $p^{\mathrm{jet}}_{T}$ cut is nearly imperceptible. This indicates that the modifications observed in the full matter$+$lbt results are completely governed by the medium effect in the low-virtuality lbt phase.
In the case of quark jets, it is conceivable that hard splittings induced by the medium at low virtuality can have larger angles than those driven by virtuality. This might be attributed to the relatively lower initial virtuality in quark jets, where the virtuality-driven splittings are less likely to have very large angles. Such large-angle medium-induced splittings can be identified as hard branches by the SoftDrop grooming, subsequently manifesting in the $r_{g}$ distribution as a shift from smaller $r_{g}$ to moderate $r_{g}$ ($\approx 0.1$), appearing as a bump structure in the full matter$+$lbt results. Conversely, for gluon jets, despite the presence of medium effects at low virtuality, splittings driven mainly by initial virtuality are still identified as hard branches by the Soft Drop grooming. The medium effects at low virtuality, which can involve large-angle but soft radiation, do not alter the intrinsic hard structure of gluon jets while still resulting in significant jet energy loss.
Figure 10: Same as Fig. 7 for jets from hard scatterings in $p$-$p$ collisions at $\sqrt{s}=5.02$ TeV, generated by pythiagun with ISR and MPI. The results are shown for $\gamma$-tagged jets (solid) from prompt photon-generating hard processes (HardQCD:all=off+PromptPhoton:all=on) and inclusive jets (dashed) from inclusive hard processes (HardQCD:all=on+PromptPhoton:all=on) generated at leading order by pythia 8 with different $p^{\mathrm{jet}}_{\mathrm{T}}$ triggers. For $\gamma$-tagged jets, isolation requirement, relative azimuth angle cut, and the additional cut of $p^{\mathrm{jet}}_{\mathrm{T}}<p^{\gamma}_{\mathrm{T}}$ are imposed.
@@
Figure 11: Same as Fig. 8 for jets from hard scatterings at $\sqrt{s_{NN}}=5.02$ TeV, generated by pythiagun with ISR and MPI. The results are shown for $\gamma$-tagged jets (solid) from prompt photon-generating hard processes (HardQCD:all=off+PromptPhoton:all=on) and inclusive jets (dashed) from inclusive hard processes (HardQCD:all=on+PromptPhoton:all=on) generated at leading order by pythia 8 with different $p^{\mathrm{jet}}_{\mathrm{T}}$ triggers. For $\gamma$-tagged jets, isolation requirement, relative azimuth angle cut, and the additional cut of $p^{\mathrm{jet}}_{\mathrm{T}}<p^{\gamma}_{\mathrm{T}}$ are imposed.
The $r_{g}$ distribution modifications for $\gamma$-tagged jets and inclusive jets in Pb-Pb collisions at $\sqrt{s_{NN}}=5.02$ TeV are compared in Fig. 11. Here, it should be noted that two factors contribute to the energy loss selection bias effect of $p^{\mathrm{jet}}_{T}$ cut, causing the $r_{g}$ narrowing irrelevant to the actual structural modification of hard splittings. Firstly, as confirmed by the results with a fixed $E_{\mathrm{init}}$, jets with originally wider splittings tend to lose more energy, resulting in less triggered jets as $r_{g}$ increases. In addition, as mentioned above for the $p$-$p$ case, jets with parent partons with larger initial $p_{T}$, which possess the narrower $r_{g}$ distribution, are triggered more likely in the presence of energy loss.

### v2

@@ III Analyses
This section provides a detailed description of the soft-drop-grooming procedure used for calculating observables that characterize the hard component structures of jets, as well as the method for constructing $\gamma$-tagged jets in our Monte Carlo simulations of high-energy heavy-ion collisions.
III.1 Soft-drop-grooming procedure
In this study, we focus on observables that characterize the structure formed by the parton shower and eventually captured as the hard component of reconstructed jets, as identified through the soft-drop-grooming algorithm [91]. The soft-drop-grooming identifies the hard splitting, while removing the soft branchings, at as large an angle as possible.
In the soft-drop-grooming procedure, an angular-ordered clustering tree is rebuilt using the Cambridge-Aachen (C/A) algorithm [68, 141] for constituents of a triggered jet reconstructed by a standard jet-finding algorithm, such as the anti-$k_{t}$ algorithm [42], with a jet cone size $R$. Next, we traverse back, i.e., from the branch with the largest angle, through the C/A tree. At each branching, we check whether the two prongs of the branching satisfy the soft-drop condition given by
$\frac{p_{T,2}}{p_{T,1}+p_{T,2}}>z_{\mathrm{cut}}\left(\frac{\Delta R_{12}}{R}\right)^{\beta},$ (1)
where $p_{T,1}$ and $p_{T,2}$ ($p_{T,1}$) are the transverse momenta of the prongs, and $\Delta R_{12}=[(\eta_{1}-\eta_{2})^{2}+(\phi_{1}-\phi_{2})^{2}]^{1/2}$ is the radial distance between the prongs in the rapidity-azimuthal angle plane. The soft-drop parameters $z_{\mathrm{cut}}$ and $\beta$ control the grooming procedure.
If the soft-drop condition is satisfied, we stop the procedure, and the two prongs of the branching are used to calculate the groomed jet observables. If the condition is not met, we continue traversing the tree by following the prong with the larger $p_{T}$ and repeat the same procedure. It should be noted that there may be cases where a prong pair satisfying the soft-drop condition cannot be found eventually. Throughout this work, the jet reconstruction and soft-drop grooming are performed using the fastjet package [44, 43] with fastjet-contrib-1.045 [1].
@@ III.2 $\boldsymbol{\gamma}$-tagged jet
However, experimentally, $\gamma$-jet pairs can also arise from processes other than the initial hard $\gamma$-jet pair creation, such as photon radiation in final-state parton showering. Therefore, to suppress contamination from such non-targeted processes, specific cuts are imposed in the experimental analyses: the isolation requirement for photons and the relative azimuth angle cut. To enhance quantitative accuracy in our comparison to experimental data, we impose these cuts on our simulated events.
@@
The relative azimuth angle cut
$\lvert\phi_{\mathrm{jet}}-\phi_{\gamma}\rvert>\Delta\phi_{\mathrm{cut}}$ (4)
@@ IV Results
This section systematically investigates the modification of the soft-drop substructures from simulations with simplified initial events generated by the pgun module. A single quark or gluon is generated at one point in the medium and allowed to propagate outward. Strict control over the flavor and energy of the parton allows for a detailed investigation of the flavor and energy dependence of specific substructure observables.
@@ IV.1 Jet splitting momentum fraction
where $p_{T,1}$ and $p_{T,2}$ ($p_{T,1}$) are transverse momenta of the pair prongs passing the soft-drop condition.
Figure 1: Distributions of jet splitting momentum fraction $z_{g}$ normalized by the number of all triggered jets (upper panel) and the number of jets passing the soft-drop condition (lower panel) for the leading jets in events generated with pgun. The jet shower evolution is performed by vacuum matter for the parent parton having $E_{\mathrm{init}}=140$ GeV. Jets are reconstructed with $R=0.4$ at midrapidity $|\eta_{\mathrm{jet}}|<2.0$. The results are shown for quark jets (solid) and gluon jets (dashed) with different $p^{\mathrm{jet}}_{T}$ triggers, 112, 84, 56, and 28 GeV. The soft-drop parameters are $z_{\mathrm{cut}}=0.2$ and $\beta=0$.
In Fig. 1, the results for leading jets from pgun simulations for the vacuum case are shown. The initial energy of the parent parton is fixed at $E_{\mathrm{init}}=140$ GeV. In the upper panel, the $z_{g}$ distribution normalized by the number of all triggered jets, including those that did not pass the soft-drop condition,
@@
is shown, whereas in the lower panel, the one normalized by the number of jets passing the soft-drop condition,
@@
is shown. Here, $N_{\mathrm{jet}}$ and $N_{\mathrm{SD,jet}}$ are the number of triggered jets and the number of jets passing the soft-drop condition, respectively.
Gluon jets pass the soft-drop condition more often than quark jets, and as a result, for the case of the normalization by the number of all triggered jets, the gluon jet consistently exhibits larger values than the quark jet across the entire $z_{g}$ region. However, when normalized by the number of jets passing the soft-drop condition, the difference between quark jets and gluon jets is nearly negligible.
Figure 2: Ratios of $z_{g}$ distributions for the leading jets in events with the parent parton having a fixed initial energy $E_{\mathrm{init}}=140$ GeV generated by pgun. The jet shower evolution in the QGP medium produced in 0%–10% Pb-Pb collisions at $\sqrt{s_{NN}}=5.02$ TeV is performed by matter++lbt. The results are shown for quark jets (solid) and gluon jets (dashed) with different $p^{\mathrm{jet}}_{T}$ triggers, 112, 84, 56, and 28 GeV. The soft-drop parameters are $z_{\mathrm{cut}}=0.2$ and $\beta=0$.
@@
Figure 3: Same as Fig. 2 for matter++lbt simulations in which the medium effects from recoils and holes are artificially removed.
Next, we explore the impact of the medium response effects via recoils and holes. Figure 3 shows the results of the pgun simulations with matter++lbt, in which the effects of medium response are artificially excluded by discarding recoil and hole partons without passing them to the hadronization module. In this case, for both quark and gluon jets, the modification of the $z_{g}$ distribution is nearly invisible, indicating that the modifications observed in the full matter++lbt results are primarily governed by medium response: recoils tend to populate the smaller momentum-fraction branch of the prong pair identified as the hard branching by soft drop. This is consistent with findings reported in studies employing jewel [104], a Monte Carlo model for in-medium jet shower evolution that includes medium response effects from recoils.
@@
Then, we delve into the detailed exploration of how the medium effects at high virtuality and low virtuality, respectively, bring about the hard splitting modification. To achieve this, we conduct simulations incorporating only the medium effects of matter, by turning off the medium effects at low virtuality ($Q<2$ GeV), and present the results of these simulations with the pgun initial hard process in Fig. 4. Note that in this configuration, recoils and holes from the matter phase are included. In this case, no medium effects are seen for either quark or gluon jets, and the recoils contributing to the soft-drop-reconstructed hard branching arise mainly in the low-virtuality phase, as modeled by the lbt module.
Figure 5: Same as Fig. 1 for jets from hard scatterings in $p$-$p$ collisions at $\sqrt{s}=5.02$ TeV, generated by pythiagun with ISR and MPI. The results are shown for $\gamma$-tagged jets (solid) from prompt photon-generating hard processes (HardQCD:all=off+PromptPhoton:all=on) and inclusive jets (dashed) from inclusive hard processes (HardQCD:all=on+PromptPhoton:all=on) generated at leading order by pythia 8 and with different $p^{\mathrm{jet}}_{\mathrm{T}}$ triggers. For $\gamma$-tagged jets, the isolation requirement, the relative azimuth angle cut, and the additional cut of $p^{\mathrm{jet}}_{\mathrm{T}}<p^{\gamma}_{\mathrm{T}}$ are imposed.
@@
For the inclusive jet analysis, events with initial scatterings, including all hard QCD 2-to-2 events (HardQCD:all=on+PromptPhoton:all=on) are used. In the distributions normalized by the number of all triggered jets (upper panel), it can be observed that inclusive jets are slightly more likely to pass the soft-drop condition compared with $\gamma$-triggered jets, due to the larger fraction of gluon jets. When the normalization is done by the number of jets passing the soft-drop condition, the difference between $\gamma$-triggered jets and inclusive jets disappear, similar to the erasing of differences between gluon and quark jets as seen in the pgun results.
Figure 6: Same as Fig. 2 for jets from hard scatterings at $\sqrt{s_{NN}}=5.02$ TeV, generated by pythiagun with ISR and MPI. The results are shown for $\gamma$-tagged jets (solid) from prompt photon-generating hard processes (HardQCD:all=off+PromptPhoton:all=on) and inclusive jets (dashed) from inclusive hard processes (HardQCD:all=on+PromptPhoton:all=on) generated at leading order by pythia 8 with different $p^{\mathrm{jet}}_{\mathrm{T}}$ triggers. For $\gamma$-tagged jets, the isolation requirement, the relative azimuth angle cut, and the additional cut of $p^{\mathrm{jet}}_{\mathrm{T}}<p^{\gamma}_{\mathrm{T}}$ are imposed.
Figure 6 shows the medium modification for the $\gamma$-tagged and inclusive jets in Pb-Pb collisions at $\sqrt{s_{NN}}=5.02$ TeV with the same triggers. For both $\gamma$-tagged jets and inclusive jets, modification becomes less significant as $p^{\mathrm{jet}}_{T}$ increases. Especially for jets with $p^{\mathrm{jet}}_{T}>300$ GeV, almost no changes in the shape of the $z_{g}$ distribution are observed. In all $z_{g}$ regions, jets passing the soft-drop condition are almost uniformly reduced due to medium effects, with the reduction being slightly more pronounced for inclusive jets.
Figure 7: Same as Fig. 2 for $\gamma$-tagged jets from prompt photon-generating hard processes (HardQCD:all=off+PromptPhoton:all=on) generated at leading order by pythia 8 at $\sqrt{s_{NN}}=5.02$ TeV for different $x_{J\gamma}$ ranges. The photon of $140<p^{\gamma}_{\mathrm{T}}<160$ GeV is triggered with the isolation requirement, and the relative azimuth angle cut.
Figure 7 shows our prediction for the $x_{J\gamma}$ dependence on the modification of the $z_{g}$ distribution for $\gamma$-tagged jets in Pb-Pb collisions at $\sqrt{s_{NN}}=5.02$ TeV. As seen in the pgun results, there is no significant $p^{\mathrm{jet}}_{T}$-selection bias effect on the $z_{g}$ distribution, and thus, the modification remains almost consistent across all presented $x_{J\gamma}$ ranges.
@@ IV.2 Jet splitting radius
where $\eta_{1}$, $\eta_{2}$, $\phi_{1}$, and $\phi_{2}$ are the rapidities and azimuthal angles of the pair prongs passing the soft-drop condition.
Figure 8: Distributions of jet splitting momentum radius $r_{g}$ normalized by the number of all triggered jets for the leading jets in events generated with pgun. The jet shower evolution is performed by vacuum matter for the parent parton having $E_{\mathrm{init}}=140$ GeV. Jets are reconstructed with $R=0.4$ at midrapidity $|\eta_{\mathrm{jet}}|<2.0$. The results are shown for quark jets (solid) and gluon jets (dashed) with different $p^{\mathrm{jet}}_{T}$ triggers, 112, 84, 56, and 28 GeV. The soft-drop parameters are $z_{\mathrm{cut}}=0.2$ and $\beta=0$.
@@
for leading gluon jets and quark jets from the vacuum pgun simulations are compared. The gluon jets exhibit much wider distributions with a peak at larger $r_{g}$ values than quark jets. This is due to the fact that gluon jets, having a larger Casimir factor and radiating more, tend to be produced with a larger virtuality compared with quark jets. For the jet cone size $R=0.4$, almost no $p^{\mathrm{jet}}_{T}$ cut dependence can be observed in the vacuum case with fixed $E_{\mathrm{init}}=140$ GeV.
Figure 9: Ratios of $r_{g}$ distributions for the leading jets in events with the parent parton having a fixed initial energy $E_{\mathrm{init}}=140$ GeV generated by pgun. The jet shower evolution in the QGP medium produced in 0%–10% Pb-Pb collisions at $\sqrt{s_{NN}}=5.02$ TeV is performed by matter++lbt. All the setups are the same as in Fig. 2.
Figure 9 shows the modification of the $r_{g}$ distribution for the pgun jets by the medium created in central Pb-Pb collisions at $\sqrt{s_{NN}}=5.02$ TeV. For gluon jets, a monotonically decreasing trend with increasing $r_{g}$ is observed in cases with a larger $p^{\mathrm{jet}}_{T}$ cut. Then, as the $p^{\mathrm{jet}}_{T}$ cut value is reduced, the decrease rate becomes more gentle, and almost no modification can be seen for $p^{\mathrm{jet}}_{T}/E_{\mathrm{init}}>0.2$. Thus, for gluon jets, the modification pattern is entirely brought by the effect from jet selection with $p^{\mathrm{jet}}_{T}$ cut: Jets with an originally large splitting radius tend to have their constituents reach the edge of the jet cone more easily, making them lose more energy and thus less likely to be triggered.
The selection bias effect with $p^{\mathrm{jet}}_{T}$ cut can be observed also for quark jets as large-$r_{g}$ suppression, with the same trend disappearing as the $p^{\mathrm{jet}}_{T}$ cut value is reduced. Additionally, in the case of quark jets, a prominent bump structure accompanied by slight suppression at small $r_{g}$ is observed in the mid-$r_{g}$ region. This feature is significantly different from that of gluon jets, providing evidence of direct modifications to the hard-branching structure within quark jets.
Figure 10: Same as Fig. 9 for matter++lbt simulations in which the medium effects from recoils and holes are artificially removed.
In Fig. 10, we show the matter++lbt results with the medium response effects from recoils and holes artificially disabled. For gluon jets, small recoil contributions to hard branchings at narrow angles manifest as a minor suppression at very small $r_{g}$ relative to the full results. In quark jets, the pronounced bump seen with recoils and holes is largely diminished, demonstrating that the feature is primarily driven by medium response, although a slight broadening remains visible. Since the recoils and holes are discarded, explicitly violating energy-momentum conservation, the large-$r_{g}$ suppression does not disappear even when the $p^{\mathrm{jet}}_{T}$ cut value is reduced.
@@
Figure 11 shows the matter-only results, in which the medium effect is turned off for $Q<2$ GeV. For both quark jets and gluon jets, $r_{g}$ distribution modification for any $p^{\mathrm{jet}}_{T}$ cut is nearly imperceptible. This indicates that the modifications observed in the full matter++lbt results are completely governed by the medium effect in the low-virtuality lbt phase.
In the case of quark jets, it is conceivable that soft-drop-reconstructed hard branchings modified by medium effects at low virtuality, primarily through the contribution of recoils, occur at larger angles than radiations predominantly driven by virtuality. This might be attributed to the relatively lower initial virtuality in quark jets, where the virtuality-driven splittings are less likely to have very large angles. Such large-angle medium-induced branchings can be identified as hard branchings by the soft-drop grooming, subsequently manifesting in the $r_{g}$ distribution as a shift from smaller $r_{g}$ to moderate $r_{g}$ ($\approx 0.1$), appearing as a bump structure in the full matter++lbt results. Conversely, for gluon jets, despite medium effects at low virtuality, splittings primarily driven by the initial virtuality are still predominantly identified as hard branchings by the soft-drop grooming. The medium effects at low virtuality do not alter the intrinsic hard structure of gluon jets while still resulting in significant jet energy loss.
Figure 12: Same as Fig. 8 for jets from hard scatterings in $p$-$p$ collisions at $\sqrt{s}=5.02$ TeV, generated by pythiagun with ISR and MPI. The results are shown for $\gamma$-tagged jets (solid) from prompt photon-generating hard processes (HardQCD:all=off+PromptPhoton:all=on) and inclusive jets (dashed) from inclusive hard processes (HardQCD:all=on+PromptPhoton:all=on) generated at leading order by pythia 8 with different $p^{\mathrm{jet}}_{\mathrm{T}}$ triggers. For $\gamma$-tagged jets, the isolation requirement, the relative azimuth angle cut, and the additional cut of $p^{\mathrm{jet}}_{\mathrm{T}}<p^{\gamma}_{\mathrm{T}}$ are imposed.
@@
Figure 13: Same as Fig. 9 for jets from hard scatterings at $\sqrt{s_{NN}}=5.02$ TeV, generated by pythiagun with ISR and MPI. The results are shown for $\gamma$-tagged jets (solid) from prompt photon-generating hard processes (HardQCD:all=off+PromptPhoton:all=on) and inclusive jets (dashed) from inclusive hard processes (HardQCD:all=on+PromptPhoton:all=on) generated at leading order by pythia 8 with different $p^{\mathrm{jet}}_{\mathrm{T}}$ triggers. For $\gamma$-tagged jets, the isolation requirement, the relative azimuth angle cut, and the additional cut of $p^{\mathrm{jet}}_{\mathrm{T}}<p^{\gamma}_{\mathrm{T}}$ are imposed.
The $r_{g}$ distribution modifications for $\gamma$-tagged jets and inclusive jets in Pb-Pb collisions at $\sqrt{s_{NN}}=5.02$ TeV are compared in Fig. 13. Here, it should be noted that two factors contribute to the energy loss selection bias effect of $p^{\mathrm{jet}}_{T}$ cut, causing the $r_{g}$ narrowing irrelevant to the actual structural modification of hard splittings. First, as confirmed by the results with a fixed $E_{\mathrm{init}}$, jets with originally wider splittings tend to lose more energy, resulting in less triggered jets as $r_{g}$ increases. In addition, as mentioned above for the $p$-$p$ case, jets with parent partons with larger initial $p_{T}$, which possess the narrower $r_{g}$ distribution, are triggered more likely in the presence of energy loss.